OpenStudy (anonymous):

solve the following eqn. \[ e^x-e^{-x}=0\] for x

hartnn (hartnn):

e^x=e^(-x) comparing the exponents x=-x

hartnn (hartnn):

can u find x from x=-x ?

hartnn (hartnn):

hint, add x on both sides.

OpenStudy (anonymous):

so its 2x=0 ?

OpenStudy (anonymous):

:S

hartnn (hartnn):

yes, x=0 is correct :)

OpenStudy (anonymous):

oh. is that all i do? :S

hartnn (hartnn):

yes, thats it 2x=0, so x=0 is the solution.

OpenStudy (anonymous):

thankyou

hartnn (hartnn):

welcome ^_^